OpenStudy (anonymous):

Please help. Write an equation for the line that is parallel to the given line and that passes through the given point. y=3/4x-9; (-8,-18) A. y=-4/3x+12 B. y=4/3x-12 C. y=3/4x+11/2 D. y=3/4x-12

OpenStudy (anonymous):

What would be the slope an equation that is parallel to?\[y = \frac{ 3 }{ 4 }x - 9\]

OpenStudy (anonymous):

The slope would be the same

OpenStudy (anonymous):

Because you already have the slope and a point, you can substitute the values of y=mx+b, with the slope and y and x values given. Therefore, y=mx+b -18=(3/4)(-8)+b -18=-6+b +6 +6 ----------- -12=b

OpenStudy (anonymous):

That would mean the equation is \[y=\frac{ 3 }{ 4 }x-12\]

OpenStudy (anonymous):

The answer is D!